# Exploring lattice thermal conductivity models via interpretable deep learning to accelerate the discovery of novel materials

### v2

@@ II.2 Deepening physical interpretability through sensitivity analysis
Figure 3: Feature correlation and sensitivity analysis. (a) A heatmap of correlation coefficients is presented, where the numbers in the upper-right triangle represent the Pearson correlation coefficients between pairs of features. In the lower-left triangle, the direction of the ellipse’s long axis indicates the positive or negative correlation, while the area reflects the strength of correlation. (b) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
@@
As shown in Figure 3, we calculated the mean absolute values of various sensitivity indices. The first point to note is that, regardless of the machine learning model used, the results obtained under the assumption of feature independence are not convincing. In Figure 3(b), the results from different methods vary noticeably, and it is difficult to determine which is more reliable. In Figure 3(d), although the trends from different analysis methods are similar, the related features should have similar sensitivity[65], a conclusion that is clearly contradicted by Figure 3(b) and (d). Therefore, the conclusions based on traditional sensitivity analysis methods lack credibility. The sensitivity analysis results in Figure 3(c) and (e) is somewhat more reasonable, but differences still exist. For example, in Figure 3(e), the importance of heat capacity appears to be underestimated. Moreover, when $\Theta$ is removed, despite the presence of other highly linearly correlated physical quantities, the performance of the SISSO model deteriorates significantly. This suggests that SISSO fails to capture the correlations between physical quantities sensitively, whereas the KAN model can. A ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ denotes a robust correlation with other features, while ${{K}^{\text{T}}}=0$ denotes an exact correlation with the other inputs[28, 66]. From this perspective, feature dependence remains a notable limitation in the SISSO model.
@@
Ultimately, we retained ${{C}_{\text{v}}}$, $\alpha$, $F$, and ${{K}_{\text{s}}}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while being conducive to prediction, facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of LTC with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. Unlike KAN, SISSO restricts each descriptor to a maximum of 7 features, which limits its expressive capacity and prevents it from capturing complex physical feature-mapping relationships.
After key feature extraction, we performed symbolic regression again using KAN and SISSO. The analytical expression fitted by KAN became considerably simplified, as shown below:
@@ II.3 Screening potential thermal insulators/conductors
Figure 5: Correlation between the 4 features and LTC. (a)(̃d) The two-dimensional mapping of different features with $\log(\kappa)$, where the color bar reflects the scatter density, and the feature values have been min-max normalized. (e) The mapping between $\log(\kappa)$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa)$ of each sample. The dashed line represents the mean $\log(\kappa)$ value of 1 for the samples.
@@ IV.4 Kucherenko Indices
Unlike metrics such as Sobol, LIME, and SHAP, the Kucherenko indices accounts for the dependencies among input features. As an extension of the Sobol sensitivity indices, the Kucherenko indices is specifically designed to quantify the sensitivity of model outputs to input variables while considering these interdependencies[63, 66, 82]. The Kucherenko indices essentially involves computing Sobol indices after first employing Copulas[83] (typically Gaussian Copulas) to separate each feature’s marginal distribution from the dependency structure[84]. This approach approximates the correlated variables as independent before calculating the Sobol indices.
@@ Acknowledgements.
We would like to acknowledge the National Key R&D Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No. 1212200482, No. 12404062, No.12302220, No.1237041184), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO. ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No. 2042023kf0109). We sincerely thank the High-Preformance Computing Center of Wuhan University for providing computational resources for our work.

### v3

@@ II.2 Deepening physical interpretability through sensitivity analysis
Figure 3: Feature correlation and sensitivity analysis. (a) A heatmap of correlation coefficients is presented, where the numbers in the upper-right triangle represent the Pearson correlation coefficients between pairs of features. In the lower-left triangle, the direction of the ellipse’s long axis indicates the positive or negative correlation, while the area reflects the strength of correlation. (b)-(e) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
@@
As shown in Figure 3, we calculated the mean absolute values of various sensitivity indices, which provide the global values for SHAP and LIME. It is important to note that, regardless of the white-box machine learning surrogate model used, results obtained under the assumption of feature independence lack persuasiveness. In Figure 3(b), significant differences exist among the results from different methods, making it difficult to determine which is more reliable. In Figure 3(d), although different analysis methods yield similar feature sensitivities, correlated features should exhibit similar sensitivities[65]. Evidently, the conclusions from Figure 3(b) and (d) do not satisfy this premise. Therefore, conclusions drawn from sensitivity analysis methods that ignore feature interaction effects lack credibility.
The sensitivity analysis results in Figure 3(c) and (e) are relatively reasonable and satisfy the premise of ”similar sensitivities for correlated features,” but differences still exist. For instance, in Figure 3(e), the total effect indices of $C_{p}$ and $C_{v}$ are higher than those of other features, yet their first-order effect indices are approximately zero. A total effect index slightly higher than the first-order effect index indicates that the independent contributions of the features themselves are minimal, but their combinations with other features exert additional influence on the model output. In the symbolic SISSO model (Eq. 1), specific heat capacities appear in the form $\left|{{C_{v}}-{C_{p}}}\right|$. Such interaction terms reduce the contribution of individual specific heat variables but result in a higher total contribution. However, a first-order effect index of zero suggests that log($\kappa$) is entirely independent of specific heat capacities.
When considering empirical models like Slack’s, the sensitivity analysis results of the SISSO symbolic model are contradictory and unsustainable, as they underestimate the influence of specific heat capacities on $log(\kappa)$. Furthermore, despite the presence of other highly correlated variables, the removal of redundant features causes a noticeable performance degradation in the SISSO model, as shown in Figure 4(a). This indicates that SISSO fails to accurately capture the correlations between physical quantities and exhibits poor robustness, whereas the KAN model shows the opposite behavior. For a primary feature, a ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ indicates a strong correlation with other features, while ${{K}^{\text{T}}}=0$ signifies an exact correlation with other inputs[28, 66]. From this perspective, KAN is more effective than SISSO at overcoming feature dependency. This advantage can be attributed to KAN’s neural network architecture. As illustrated in Figure 4(b), in the absence of pruning, the neurons in KAN’s hidden layers are fully connected to both preceding and subsequent layers. This design allows each feature to independently and equitably contribute to log($\kappa$). Simultaneously, KAN leverages backpropagation to update the weights of various nodes, ensuring the model achieves optimal accuracy. This process not only guarantees fair participation of all features in model training but also effectively distinguishes their relative contribution.
@@
Ultimately, we retained ${{C}{\text{v}}}$, $\alpha$, $F$, and ${{K}{\text{s}}}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of LTC with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq.13), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
After key feature extraction, we performed symbolic regression again using KAN and SISSO. The symbolic model fitted by KAN became considerably simplified, as shown below:
@@ II.3 Screening potential thermal insulators/conductors
Figure 5: Correlation between the 4 features and LTC. (a)-(d) The two-dimensional mapping of different features with $\log(\kappa)$, where the color bar reflects the scatter density, and the feature values have been min-max normalized. (e) The mapping between $\log(\kappa)$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa)$ of each sample. The dashed line represents the mean $\log(\kappa)$ value of 1 for the samples.
@@ IV.4 Kucherenko Indices
Unlike metrics such as Sobol, LIME, and SHAP, the Kucherenko indices accounts for the dependencies among input features. As an extension of the Sobol sensitivity indices, the Kucherenko indices is specifically designed to quantify the sensitivity of model outputs to input variables while considering these interdependencies[63, 66, 82]. The Kucherenko indices essentially involves computing Sobol indices after first employing Copulas[83] (typically Gaussian Copulas) to separate each feature’s marginal distribution from the dependency structure[84], as detailed in Table S2. This approach approximates the correlated variables as independent before calculating the Sobol indices. The Gaussian copula is constructed based on the standard normal distribution, and its mathematical expression is:
$C\left({{u_{1}},{u_{2}},\ldots,{u_{d}};\Sigma}\right)={\Phi_{\Sigma}}\left[{{\Phi^{-1}}\left({{u_{1}}}\right),{\Phi^{-1}}\left({{u_{2}}}\right)\ldots,{\Phi^{-1}}\left({{u_{d}}}\right)}\right]$ (18)
where $u_{i}=F_{i}(x_{i})$ is the Cumulative Distribution Function (CDF) of the marginal distribution, $\Phi^{-1}$ is the inverse of the CDF of the standard normal distribution $\mathcal{N}(0,1)$, $\Phi_{\Sigma}$ is the CDF of a multivariate normal distribution with mean zero and covariance matrix $\Sigma$, and $\Sigma$ is the correlation matrix, describing the dependence structure between the random variables.
@@ Acknowledgements.
We would like to acknowledge the National Key R&D Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No.12122408, No.12404062, No.12302220, No.1237041184), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO. ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No. 2042023kf0109). We sincerely thank the High-Preformance Computing Center of Wuhan University for providing computational resources for our work.
Supplementary Information
Supplementary Note 1 KAN’s symbolic model on complete dataset
The simplified symbolic model in the main text Eq.(5), derived from key physical features, is sufficient for an accurate description of LTC. However, for reference, we also provide the symbolic expression fitted by KAN using the complete dataset. It is worth noting that, due to the architecture of KAN, this formula is inevitably more complex. Therefore, we present it solely in its Python representation:
1 log(kappa) =
[…]
Supplementary Note 2 Multi-Layer Perceptrons (MLPs)
The theoretical framework of MLPs is well-established. Therefore, in this work, we emphasize the distinctions between MLPs and KANs. Unlike KANs, MLPs are founded on the universal approximation theorem[100], which can be mathematically expressed as:
$f\left(x\right)\approx\sum\limits_{i=1}^{N}{{{\alpha}_{i}}\sigma\left(\omega_{i}^{\text{T}}x+{{b}_{i}}\right)}$ (Supplementary Eq.1)
where $\alpha$ are the coefficients, $\sigma$ is the activation function (we adopted ReLU[101] in this work), ${{\omega}_{i}}$ and ${{b}_{i}}$ denote the weights and biases, $N$ represents the number of neurons in the hidden layer. The expression of MLPs can be simplified to:
$\text{MLP}\left(\mathbf{x}\right)=\left({{\mathbf{W}}_{L-1}}\circ\sigma\circ{{\mathbf{W}}_{L-1}}\circ\sigma\circ\cdots\circ{{\mathbf{W}}_{1}}\circ\sigma\circ{{\mathbf{W}}_{0}}\right)\mathbf{x}$ (Supplementary Eq.2)
In MLPs, only the weights and biases, which serve as linear parameters, are learnable, while the nonlinear activation functions are fixed as hyperparameters.
Supplementary Note 3 eXtreme Gradient Boost (XGB)
XGB represents a great advancement over the traditional Gradient Boosting Decision Tree (GBDT) algorithm. Its innovations include the incorporation of regularization to enhance generalization, a two-step gradient approximation of the objective function to accelerate computational efficiency, column subsampling to mitigate noise and further bolster generalization, and an advanced method for handling missing values by assigning a default classification direction at tree nodes, which renders it particularly effective for sparse datasets[102]. GBDT algorithms, including XGB, are computationally intensive particularly in tasks with high-dimensional features and large datasets[102, 103].
Supplementary Note 4 Crystal Graph Convolutional Neural Network (CGCNN)
CGCNN[104] directly learns from the connection of atoms in a crystal, representing the crystal as a graph with atoms as nodes and bonds as edges. This graph-based approach enables the model to automatically extract optimal representations for predicting material properties. The key feature of CGCNN is its Convolutional Neural Network (CNN) built on top of these crystal graphs. The convolutional layers update the feature vectors of atoms by incorporating information from neighboring atoms and bonds, using the following graph convolution function:
$v_{i}^{\left(t+1\right)}=\text{Conv}\left(v_{i}^{\left(t\right)},v_{j}^{\left(t\right)},u_{\left(i,j\right)}^{\left(k\right)}\right),{{\left(i,j\right)}_{k}}\in G$ (Supplementary Eq.3)
where $v_{i}^{\left(t\right)}$ is the feature vector of atom i at layer t, and $u_{\left(i,j\right)}^{\left(k\right)}$ is the feature vector of the bond between atoms i and j. After multiple convolutional layers, the pooling layers aggregate the atomic feature vectors into an overall crystal representation, which is then passed to fully connected layers to predict material properties.
Supplementary Note 5 SHAP Indices
Shapley Additive Explanations (SHAP)[105] is derived from the concept of Shapley values[106] in game theory, which were originally employed to assess the contribution of each individual player to the collective gain within a coalition[107]. Shapley value for individual player can be expressed as:
${{\phi}_{i}}=\sum\limits_{S\subseteq N\backslash\left\{i\right\}}{\frac{\left|S\right|!\left(\left|N\right|-\left|S\right|-1\right)!}{\left|N\right|!}}\left(v\left(S\cup\left\{i\right\}\right)-v\left(s\right)\right),i=1,2,\ldots,N$ (Supplementary Eq.4)
where $N$ represents the set of all features, $S$ is any subset of features that does not include feature $i$, $\left|S\right|$ is the number of features in set $S$, is the contribution of feature set $S$ to the model’s prediction output, and
$v\left(S\cup\left\{i\right\}\right)$
is the contribution of the feature set
$S\cup\left\{i\right\}$
, which includes feature $i$, to the model’s prediction output. The core concept of SHAP is to generate explanations for individual predictions, which can then be aggregated to provide both local and global insights into the model’s behavior[105], it can be expressed mathematically as:
$\hat{f}\left(\mathbf{x}\right)={{\phi}_{0}}+\sum\limits_{i=1}^{N}{{{\phi}_{i}}\left(\mathbf{x}\right)}$ (Supplementary Eq.5)
where
${{\phi}_{i}}\left(\mathbf{x}\right)$
is the Shapley value of the i-th variable, ${{\phi}_{0}}$ is typically set as the expected value of the training samples, $\hat{f}\left(\mathbf{x}\right)$ represents the original prediction made by the model[108].
Supplementary Note 6 LIME Indices
Local Interpretable Model-agnostic Explanations (LIME)[60] is a perturbation-based method for feature importance analysis that employs a local surrogate model to interpret individual samples. By generating perturbations around the selected sample, a new dataset is created, and an interpretable model (such as linear regression) is trained. The difference between the two models is then measured using an objective function:
$\xi\left(x\right)=\arg\underset{g\in G}{\mathop{\min}}\,L\left(f,g,{{\pi}_{x}}\right)+\Omega\left(g\right)$ (Supplementary Eq.6)
in the expression, $f$ denotes the model to be explained, $g$ represents a simplified model drawn from a set $G$ of potential models (such as linear models), ${{\pi}_{x}}$ quantifies the distance between new samples ${x}^{\prime}$ and the original samples $x$, and $\Omega\left(g\right)$ reflects the complexity of the model $g$.
Supplementary Note 7 Supplementary Tables
Supplementary Table 1: The optimal hyperparameters.
Model Hyperparameter Optimal Value
MLP hidden_sizes [64, 64, 64] (fixed)
learning_rate 0.002587427
batch_size 186
num_epochs 592
XGB n_estimators 489
learning_rate 0.134952344
max_depth 6
subsample 0.824230319
colsample_bytree 0.869698424
reg_alpha 0.9
reg_lambda 37.30538969
KAN width 7
grid 15
k 5
lamb_l1 9.559785284
steps 27
lr 0.47852298
SISSO (10-D) desc_dim 2
nsf 10
ops (+)(-)(*)(/)(exp)(exp-)(^-1)(^2)(^3)(sqrt)(cbrt)(log)(—-—)(scd)(^6)(sin)(cos)
fcomplexity 5
funit (1:10)
nmodels 100
nf_sis 10
SISSO (4-D) desc_dim 2
nsf 4
ops (+)(-)(*)(/)(exp)(exp-)(^-1)(^2)(^3)(sqrt)(cbrt)(log)(—-—)(scd)(^6)(sin)(cos)
fcomplexity 4
funit (1:4)
nmodels 100
nf_sis 10
Supplementary Table 2: Marginal distribution of individual primary features determined by Copula sampling.
Distribution Parameters
$\Theta$ lognorm $\mu=-1.935,\sigma=0.519$
$\gamma$ lognorm $\mu=1.671,\sigma=0.012$
$C_{p}$ lognorm $\mu=-2.611,\sigma=0.778$
$C_{v}$ lognorm $\mu=-2.674,\sigma=0.831$
$\alpha$ lognorm $\mu=-2.370,\sigma=0.507$
$S_{v}$ norm $\mu=0.323,\sigma=0.119$
$F$ lognorm $\mu=-0.027,\sigma=0.102$
$K_{s}$ gamma $k=2.546,\theta=0.102$
$K_{i}$ gamma $k=2.441,\theta=0.104$
Supplementary Note 8 Supplementary Figure
Supplementary Figure 1: Histogram of the marginal distribution of a single key variable.
The variables involved in Table S2 are classified into three types of distributions: Gamma distribution, normal distribution, and log-normal distribution. Their respective probability density functions (PDF) can be expressed as:
$f\left(x\right)$ $=\frac{1}{{\Gamma\left(k\right){\theta^{k}}}}{x^{k-1}}{{\rm{e}}^{-\frac{x}{\theta}}},$ (Supplementary Eq.7)
$f(x)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right),$ (Supplementary Eq.8)
$f\left(x\right)$ $=\frac{1}{{x\sigma\sqrt{2\pi}}}{{\rm{e}}^{-\frac{{{{\left[{\ln\left(x\right)-\mu}\right]}^{2}}}}{{2{\sigma^{2}}}}}}.$ (Supplementary Eq.9)